what are all solutions of 4(x-1)^2-3=25
Answer by stanbon(75887)

4(x-1)^2-3=25
-------
4(x-1)^2 = 28
-------
(x-1)^2 = 7
-----
x-1 = sqrt(7) or x-1 = -sqrt(7)
x = 1 + sqrt(7) or x = 1-sqrt(7)
=================
Cheers,
Stan H.
=================